# Viktar Kireyeu

### v3

@@ Abstract
A new open-source cluster finding library ”Common Clusterization Library” (CCL) is proposed to describe the clusters production when applied to the transport codes. The new library was applied to the Parton-Hadron-Quantum-Molecular Dynamics approach to be comparable with the established and well known MST and SACA implementations on the exactly same set of the events. The presented procedures were tested with rapidity density distributions and complex observables like $\langle M_{imf}\rangle$ and $\langle Z_{\text{max}}\rangle$ vs $Z_{bound}2$ at the energies of ALADiN and NA49 experiments. A modification to the simulated annealing chain is proposed to reduce the computational time. The new stable clusters tracking algorithm for the ”lost” cluster recovering is presented. The results of the ”box-like” coalescence procedure and a mixed method which incorporates the coalescence algorithm steps and the negative binding energy criteria are presented. The difference between the Helium-3 found by the MST-based cluster finding procedure and the coalescence algorithm applied on the exactly same set of events and baryons is shown for the first time.
@@ 1 Introduction
Decades ago transport approaches were proposed to address the problem of the heavy-ion collisions dynamics description. It is experimentally known, that in such collisions initial nuclei can fragment into new free particles and many heavier bound objects: nuclei, hypernuclei. One of the very first attempts to describe the cluster production was the application of the cluster finding procedure in [Aichelin:1991xy], which was based on the proximity in the coordinate space criterion – so called ”Minimum Spanning Tree”, ”MST” procedure [zbMATH02560699].
Then, a new method for the earlier clusters recognition at earlier time steps, when particles (and bound groups of particles) are not yet well separated was proposed in [Dorso:1992ch] where authors firstly introduced the Simulated Annealing (SA) procedure [SA] into the transport approaches. The proposed method used the overall system binding energy as the minimization criteria for the Metropolis algorithm [METROPOLIS], but was presented for the relatively small systems.
The next evolution of the application of the Simulated Annealing technique was proposed in [Puri:1998te] for heavier initial nuclei (larger systems): the authors used the classical SA procedure as the first ”Metropolis step”, then a second ”Fragment exchange” step was applied. On this step fragments or nucleons obtained after the first step were passed again through the simulated annealing procedure to overcome the trapping to any local minima.
Recently, a new approach for the early clusters identification in the transport approaches, FRIGA, was described in [LeFevre:2019wuj]. It was demonstrated that inclusion of additional components into the binding energy calculations yields into the better description of the isotopes yields and hypernuclei production.
All proposed approaches shows good results and agrees with the experimental data in wide energy range, but all of them are closed-source and available mostly for a very limited set of the transport codes.
On order to overcome this limitation a new open-source library ”Common Clusterization Library” (CCL) [ccl_github] is developed on the base of previously presented ”psMST” [Kireyeu:2021igi] library. The CCL library is distributed under the terms of the ”Let’s Be Scientific” (”LBS”) license which allows for the free code usage and modification, but strictly requires the explicit permission by the authors prior any kind of results publication to avoid the inappropriate use of the code. The full license text is distributed with the library code.
@@
The coalescence procedure and the proposed ”mixed” procedure, which combines the coalescence algorithm steps and the usage of the $E_{\text{bind}}$ criterion, will be shown with the NA49 experimental data to be comparable with previous coalescence results within ”psMST” [Kireyeu:2023spj]. Only QMD-based approaches will be used as ”parent” models to simulate the data as in contrast to BUU-based models they can propagate particles correlations through the collision evolution without additional modifications [LeFevre:2019wuj].
@@ 2 MST, DSU
A spanning tree is a simplified version of a connected graph: it is a subgraph that uses the original graph’s vertices but only enough edges to connect every vertex without cycles. A single graph can have many different spanning trees. When each edge in the original graph has a ”cost” or ”weight”, we can look for the spanning tree with the smallest total ’weight” – this is called a Minimum Spanning Tree (MST).
@@
Usually, MST in the context of the ”clustering” in the heavy-ion collisions is understood as an algorithm which relies on the coordinate space only, however, in the authors opinion, it is not limited to this and could therefore be extended by including the momentum space too or any other physically justified criteria. Thus, a possibility to check the proximity in the momentum space between two baryons in the center of mass frame of the checked pair is also available: $\Delta p=\mathbf{p}_{i}^{\prime}-\mathbf{p}_{j}^{\prime}$, where $\mathbf{p}_{i}^{\prime}$, $\mathbf{p}_{j}^{\prime}$ are baryons momentum are in the pair COM frame.
The Disjoint Set Union (DSU), also known as Union-Find, is a data structure that efficiently manages a partition of a set of elements into disjoint (non-overlapping) subsets, for the first time described in [DSU]. It supports two primary operations:
@@
Figure 1: Upper panel: $\langle M_{imf}\rangle$ vs $Z_{bound}2$; Lower panel: $\langle Z_{\text{max}}\rangle$ vs $Z_{bound}2$; Black solid lines – PHQMD MST at $t=50$ fm/c. Red dashed lines – PHQMD MST at $t=100$ fm/c. Orange dash-dotted lines – CCL MST at $t=50$ fm/c. Blue dotted lines – CCL MST at $t=100$ fm/c. ALADiN experimental data (black circles) are taken from [LeFevre:2019wuj].
@@
Figure 2: $\langle M_{imf}\rangle$ (top panel) and $\langle E_{\text{bind}}\rangle$ (bottom) vs $Z_{bound}2$ for CCL MST. Black solid lines – $t=50$ fm/c, red dashed lines – at $t=100$ fm/c, orange dash-dotted lines – $t=150$ fm/c, blue dotted lines – $t=200$ fm/c. Black circles – ALADiN experimental data [LeFevre:2019wuj].
@@ 3 SACA, SA
The main problem with the total system binding energy check is related to the combinatorial issue. For the $Au+Au$ system with 394 baryons the number of total possible combination of free and bound particles is represented by the Bell number $B_{394}$ which is roughly $B_{394}>10^{500}$ combinations per event. Despite the fact that the Simulated Annealing procedure is limited by the control parameters such as ”temperature”, ”cooling rate”, ”number of tries” and others to speedup the calculational time, and thus can not check all available combinations, it can find the system binding energy close to the global minimum.
@@ 3.2 Full CCL SA chain
Steps 1 and 2 are the key difference between this procedure and the original SACA [Puri:1998te] idea, where the MST clusters plays the role of ’seeds’ for the first pass of the Simulated Annealing applied on the whole set of the MST clusters and free baryons.
It is more profitable in terms of the calculational time to use less initial steps for the ”pass 1” and increase it for the ”pass 2” ($\times 2$ of ”pass 1” here).
For this study the ”CCL SA” procedure was applied to the simulated data with the parameters shown in the Table 1. The number of initial steps, initial and final ”temperatures” as well as the initial probability of baryon to became ”free” are fixed for all sets, only the clusterization radius $r_{\text{clust}}$ and the momentum difference $\Delta p$ were varied.
Label $r_{\text{clust}}$ $\Delta p$ $P_{max}$ $P_{min}$ $T_{\text{max}}$ $T_{\text{min}}$ Steps
S1 2.5 0.3 0.9 0.1 1 $10^{-15}$ 500
S2 2.5 0.9 0.1 1 $10^{-15}$ 500
@@ 3.3 SACA and SA results
Figure 3: $\langle M_{imf}\rangle$ (top panel), $\langle Z_{\text{max}}\rangle$ (middle panel) and $\langle E_{bind}\rangle$ (bottom panel) as a function of $Z_{bound}2$. Black solid lines and orange dash-dotted lines are PHQMD SACA at $t=50$ and $t=100$ fm/c. Red dashed lines and blue dotted lines – CCL SA at $t=50$ and $t=100$ fm/c. Black circles – ALADiN experimental data taken from [LeFevre:2019wuj].
@@ 4 Tracking
As it was discussed in [Coci:2023daq], the QMD dynamics propagates single baryons and thus if the MST-based clusterization procedure marks group of baryons as a stable cluster, on the next time step this cluster can dissolve without collision occurred even if the corresponding quantum system must be stable. The authors proposed the ”advanced MST” (”aMST”) procedure, which uses the collision history and the information of the kinetic ”freeze-out” time to reconstruct back some of the artificially ”destroyed” clusters.
@@
After processing all $p\in\mathcal{P}$ within all subsequent time steps, the updated set $\mathcal{S}$ contains an ensemble of stable bound clusters at the end of reaction. Fig. 4 illustrates the work of the stable cluster tracking algorithm which is denoted as ”sMST” (”smart” MST) on the plots. The parameters sets used to define the MST edges are shown in the Table 2. On the two upper panels the rapidity density distributions for deuterons at the lowest NA49 experiment energy [NA49:2016qvu] are shown, on the two bottom panels same distributions are presented for the $^{3}He$ correspondingly. The (s)MST procedure with T1 set is shown with black solid lines, (s)MST with T2 set is represented by the red dashed lines, T3 set – by orange dash-dotted lines. In both cases when the tracking procedure was disabled, the clusters multiplicity in mid-rapidity region at the end of reaction are much lower than the experimental measurements. The enabling of the ”sMST” tracking procedure drastically improves the mid-rapidity clusters stability.
Figure 4: Rapidity density distribution of deuterons and $^{3}He$ with and without stable clusters tracking procedure with several parameters sets for the MST edges (2): black solid lines – T1 set, T2 – red dashed lines, T3 set – orange dash-dotted lines. Black circles – NA49 experimental data[NA49:2016qvu].
@@ 5 Coalescence
More than half of a century passed since the coalescence approach was firstly proposed [Butler:1961pr] to describe the deuterons production in emulsions, then modified to describe the light nuclei production in relativistic heavy-ion reactions [Gutbrod:1976zzr]. Now the CCL implements the ’box-like’ coalescence proposed and successfully used by the authors of the UrQMD [Bass:1998ca, Bleicher:1999xi] transport approach in [Sombun:2018yqh, Reichert:2022mek].
@@ 5.1 CCL coalescence procedure
where $P$ is a maximum momentum difference – an algorithm control parameter. The particles can coalesce if both checks are satisfied. However, since only light nuclei formation is implemented in the coalescence procedure, the QMD approaches gives a multitude of that small clusters correlated in the coordinate-momentum space, moreover, some of these smaller clusters could be parts of bigger fragments. Thus, in CCL, after finding coalescence candidates, final nuclei are chosen by the probability driven by the spin-isospin coupling (see [Sombun:2018yqh] for more details). After passing the probability cuts, the final nuclei are collected.
@@ 5.2 Coalescence results
The limited statistics of 50k central (b = 050-5 fm) $Pb+Pb$ events simulated within PHQMD transport approach at $E_{\text{lab}}=20$ AGeV was used to test the CCL coalescence procedure. Several sets of coalescence parameters ($\Delta r$, $\Delta p$) were used, all they are listed in the Table 3. The ”M1” set parameters are imported from [Reichert:2025rnw], the ”M2” parameters simulates the classical MST-based clusterization parameters, the ”M3” are taken similar to those within the ”aMST” procedure [Coci:2023daq] (increased coalescence radius) and ”M4” are the mix of all previous parameters chosen to better description of available data of the NA49 experiment at this energy ([NA49:2016qvu]).
@@
As is seen on the plot, the coalescence approach based on the ”spin-isospin coupling” not only overestimates the experimental data points, but also show the shape of the distribution with local maximum at mid-rapidity $y=0$, which differs from the experimental data shape. The amount of the ”kinetic” deuterons produced in $2\leftrightarrow 3$ reactions through the pion catalyzer is rather small at this energy, however the final result must be the sum of ”kinetic” deuterons and those found by (s)MST or formed via coalescence. The ”sMST + kinetic” sum is describes the experimental data shape quite well, however, underestimates it. There is no reason to make that similar sum for the probabilistic coalescence as it already overestimate the data.
The coalescence results (solid black line) differs from those from [Kireyeu:2023spj], but this can be explained by the smaller coalescence parameters $\Delta r=3.575$ fm and $\Delta p=0.285$ GeV/c used in previous study, which led to the smaller yield of deuterons.
@@
Figure 6: Rapidity density distribution of deuterons. Black solid lines are coalescence deuterons. The orange dash-dotted line are ”kinetic” deuterons (see text). The red dashed lines shows the sum of the ”sMST” and ”kinetic” deuterons. Black circles – NA49 experimental data (measured and mirrored) [NA49:2016qvu].
@@
Figure 7: Rapidity density distribution of $^{3}He$. Black solid lines are coalescence with spin-isospin coupling cut. The red dashed lines shows the ”sMST” results. Black circles – NA49 experimental data[NA49:2016qvu].
@@
Figure 9: Rapidity density distribution of deuterons (top panel) and $^{3}He$ (bottom) found by the coalescence based on the $E_{\text{bind}}<0$ cut. Deuteron lines are coupled with ”kinetic” deuterons. Black solid lines are for the initial ”M1” set of parameters, orange dash-dotted lines – for the ”M2” set, ”M3” set – red dashed lines and blue dotted lines are for the set ”M4”. Black circles – NA49 experimental data (measured and mirrored) [NA49:2016qvu].

### v4

@@ Abstract
A new open-source cluster finding library ”Common Clusterization Library” (CCL) is proposed to describe the clusters production when applied to the transport codes. The new library was applied to the Parton-Hadron-Quantum-Molecular Dynamics approach to be comparable with the established and well known MST and SACA implementations on the same set of the events. The presented procedures were tested with rapidity density distributions and complex observables like $\langle M_{imf}\rangle$ and $\langle Z_{\text{max}}\rangle$ vs $Z_{bound}2$ at the energies of ALADiN and NA49 experiments. A modified simulated annealing chain were used to reduce the computational time. The new ”stable” clusters tracking algorithm ”sMST” for the ”lost” QMD cluster recovering is presented. The results of the ”box-like” coalescence procedure and a mixed method which incorporates the coalescence algorithm steps and the negative binding energy criteria are presented. The difference between the Helium-3 found by the MST-based cluster finding procedure and the coalescence algorithm applied on the same set of events and baryons is shown for the first time.
@@ 1 Introduction
Decades ago transport approaches were proposed to address the problem of the heavy-ion collisions dynamics description. It is experimentally known that in such collisions initial nuclei can fragment into the new single particles and heavier bound objects: nuclei and hypernuclei. One of the very first attempts to describe the cluster production was the application of the cluster finding procedure in [1], which was based on the proximity in the coordinate space criterion – so called ”Minimum Spanning Tree”, ”MST” procedure [5].
Then, a new method for the clusters recognition at earlier time steps, when particles (and bound groups of particles) are not yet well separated was proposed in [8], where authors firstly applied the Simulated Annealing (SA) procedure [16] to the transport approaches. The proposed method used the overall system binding energy as the minimization criteria for the Metropolis algorithm [15], but was presented for the relatively small systems.
The next evolution of the SA-based technique was proposed in [17] for heavier initial nuclei (larger systems): the authors used the classical SA procedure as the first ”Metropolis step”, then a second ”Fragment exchange” step was applied. At this step fragments or nucleons obtained after the first step were passed again through the SA procedure to overcome the trapping to any local minima.
Recently, a new approach for the early clusters identification in the transport approaches, FRIGA, was described in [14]. It was demonstrated that inclusion of additional components into the binding energy calculations yields into the better description of the isotopes and hypernuclei production.
All proposed approaches shows good results and agrees with the experimental data in wide energy range, however all of them are closed-source and are available mostly for a very limited set of the transport codes if not heavily tied to the single one.
In order to overcome this limitation a new open-source library ”Common Clusterization Library” (CCL) [13] is developed on the base of previously presented ”psMST” [12] library. The CCL library is distributed under the terms of the ”Let’s Be Scientific” (”LBS”) license which allows for the free code usage and modification, but strictly requires the explicit permission by the authors prior any kind of results publication to avoid the inappropriate use of the code. The full license text is distributed with the library code.
@@
    Recovering of the lost QMD clusters during the reaction evolution.
  • 4.
@@
The coalescence and the ”mixed” procedure, which combines the coalescence algorithm steps and the usage of the $E_{\text{bind}}$ criterion, will be shown with the NA49 experimental data to be comparable with previous coalescence results within ”psMST” [11]. Only QMD-based approaches will be used as base models to simulate the data as in contrast to BUU-based models they can propagate particles correlations through the collision evolution without additional modifications [14].
@@ 2 MST, DSU
The spanning tree is a simplified version of the connected graph: it is a subgraph that uses the original graph’s vertices but only enough edges to connect every vertex without cycles. A single graph can have many different spanning trees. When each edge in the original graph has a ”cost” or ”weight”, we can look for the spanning tree with the smallest total ’weight” – this is called a Minimum Spanning Tree (MST).
@@
Usually, the ”MST” in the context of the ”clustering” in the heavy-ion collisions is understood as an algorithm which relies on the coordinate space only, however, in the authors opinion, it is not limited to this and could therefore be extended by including the momentum space too or any other physically justified criteria. Thus, a possibility to check the proximity in momentum space for baryons in the center‑of‑mass frame of the checked pair or cluster is also available. For a pair, the two baryons are boosted to their pair COM frame and the pair is accepted only if each baryon satisfies ($|\mathbf{p}_{i}^{\prime}|<p_{0}$). For a cluster, the merged system is boosted to its COM frame and the merge is accepted only if all constituents satisfy ($|\mathbf{p}_{i}^{\prime}|<p_{0}$).
The Disjoint Set Union (DSU), also known as Union-Find, is a data structure that efficiently manages a partition of a set of elements into disjoint (non-overlapping) subsets, for the first time described in [9]. It supports two primary operations:
@@
Figure 1: Upper panel: $\langle M_{imf}\rangle$ vs $Z_{bound}2$; Lower panel: $\langle Z_{\text{max}}\rangle$ vs $Z_{bound}2$; Black solid lines – PHQMD MST at $t=50$ fm/c. Red dashed lines – PHQMD MST at $t=100$ fm/c. Orange dash-dotted lines – CCL MST at $t=50$ fm/c. Blue dotted lines – CCL MST at $t=100$ fm/c. ALADiN experimental data (black circles) are taken from [14].
@@
Figure 2: $\langle M_{imf}\rangle$ (top panel) and $\langle E_{\text{bind}}\rangle$ (bottom) vs $Z_{bound}2$ for CCL MST. Black solid lines – $t=50$ fm/c, red dashed lines – at $t=100$ fm/c, orange dash-dotted lines – $t=150$ fm/c, blue dotted lines – $t=200$ fm/c. Black circles – ALADiN experimental data [14].
@@ 3 SACA, SA
The main problem with the total system binding energy check is related to the combinatorial issue. For the $Au+Au$ system with 394 baryons the number of all possible combinations of free and bound particles is represented by the Bell number $B_{394}$ which is roughly $B_{394}>10^{500}$ combinations per event. Despite the fact that the Simulated Annealing procedure is limited by the control parameters such as ”temperature”, ”cooling rate”, ”number of tries” and others to speedup the calculational time, and thus can not check all available combinations, it can find the system binding energy close to the global minimum.
@@ 3.1.4 Acceptance criterion
The early exit from the SA loop is handled by the ”stagnation” counter. Let
$L=\max\!\left(s_{\text{min}},\frac{\text{steps}}{\max(1,s_{\text{den}})}\right)$
and $c$ – counter of consecutive rejected moves. Then, for each rejected move: $c\leftarrow c+1$; for each accepted move: $c\leftarrow 0$. The simulated annealing loop is terminated early if $c\geq L$.
@@ 3.2 Full CCL SA chain
Steps 1 and 2 are the key difference between this procedure and the original SACA [17] idea, where the MST clusters plays the role of ’seeds’ for the first pass of the Simulated Annealing applied on the whole set of the MST clusters and free baryons.
It is more profitable in terms of the calculational time to use less initial steps for the ”pass 1” and increase it for the ”pass 2”. The ”stagnation”-governed early exit reduce the calculational time even more.
For this study the ”CCL SA” procedure was applied to the simulated data with the parameters shown in the Table 1.
Label $r_{\text{clust}}$ $\Delta p$ $P_{max}$ $P_{min}$ $T_{\text{max}}$ $T_{\text{min}}$ $\text{Steps}_{\text{ pass }1}$ $\text{Steps}_{\text{ pass }2}$ $s_{\text{min}}$ $s_{\text{den}}$
S1 2.5 0.3 0.9 0.1 1 $10^{-15}$ 500 4000 50 2
S2 2.5 0.9 0.1 1 $10^{-15}$ 500 4000 50 2
@@ 3.3 SACA and SA results
Figure 3: $\langle M_{imf}\rangle$ (top panel), $\langle Z_{\text{max}}\rangle$ (middle panel) and $\langle E_{bind}\rangle$ (bottom panel) as a function of $Z_{bound}2$. Black solid lines and orange dash-dotted lines are PHQMD SACA at $t=50$ and $t=100$ fm/c. Red dashed lines and blue dotted lines – CCL SA at $t=50$ and $t=100$ fm/c. Black circles – ALADiN experimental data taken from [14].
@@ 4 Tracking
As it was discussed in [7], the QMD dynamics propagates single baryons and thus if the MST-based clusterization procedure marks group of baryons as a stable cluster, on the next time step this cluster can dissolve without collision occurred even if the corresponding quantum system must be stable. The authors proposed the ”advanced MST” (”aMST”) procedure, which uses the collision history and the information of the kinetic ”freeze-out” time to reconstruct back some of the artificially ”destroyed” clusters.
@@
After processing all $p\in\mathcal{P}$ within all subsequent time steps, the updated set $\mathcal{S}$ contains an ensemble of stable bound clusters at the end of reaction. Fig. 4 illustrates the work of the stable cluster tracking algorithm which is denoted as ”sMST” (”smart” MST) on the plots. The parameters sets used to define the MST edges are shown in the Table 2. On the two upper panels the rapidity density distributions for deuterons at the lowest NA49 experiment energy [2] are shown, on the two bottom panels same distributions are presented for the $^{3}He$ correspondingly. The (s)MST procedure with T1 set is shown with black solid lines, (s)MST with T2 set is represented by the red dashed lines, T3 set – by orange dash-dotted lines. In both cases when the tracking procedure was disabled, the clusters multiplicity in mid-rapidity region at the end of reaction are much lower than the experimental measurements. The enabling of the ”sMST” tracking procedure drastically improves the mid-rapidity clusters stability.
Figure 4: Rapidity density distribution of deuterons and $^{3}He$ with and without stable clusters tracking procedure with several parameters sets for the MST edges (2): black solid lines – T1 set, T2 – red dashed lines, T3 set – orange dash-dotted lines. Black circles – NA49 experimental data[2].
@@ 5 Coalescence
More than half of a century passed since the coalescence approach was firstly proposed [6] to describe the deuterons production in emulsions, then modified to describe the light nuclei production in relativistic heavy-ion reactions [10]. The CCL implements the ’box-like’ coalescence proposed and successfully used by the authors of the UrQMD [3, 4] transport approach in [20, 19].
@@ 5.1 CCL coalescence procedure
where $P$ is a maximum momentum difference – an algorithm control parameter. The particles can coalesce if both checks are satisfied. However, since only light nuclei formation is implemented in the coalescence procedure, the QMD approaches gives a multitude of that small clusters correlated in the coordinate-momentum space, moreover, some of these smaller clusters could be parts of bigger fragments. Thus, in CCL, after finding coalescence candidates, final nuclei are chosen by the probability driven by the spin-isospin coupling (see [20] for more details). After passing the probability cuts, the final nuclei are collected.
@@ 5.2 Coalescence results
The limited statistics of 50k central (b = 050-5 fm) $Pb+Pb$ events simulated within PHQMD transport approach at $E_{\text{lab}}=20$ AGeV was used to test the CCL coalescence procedure. Several sets of coalescence parameters ($\Delta r$, $\Delta p$) were used, all they are listed in the Table 3. The ”M1” set parameters are imported from [18], the ”M2” parameters simulates the classical MST-based clusterization parameters, the ”M3” are taken similar to those within the ”aMST” procedure [7] (increased coalescence radius) and ”M4” are the mix of all previous parameters chosen to better description of available data of the NA49 experiment at this energy ([2]).
@@
As is seen on the plot, the coalescence approach based on the ”spin-isospin coupling” not only overestimates the experimental data points, but also show the shape of the distribution with local maximum at mid-rapidity $y=0$, which differs from the experimental data shape. The multiplicity of the ”kinetic” deuterons produced in $2\leftrightarrow 3$ reactions through the pion catalyzer is rather small at this energy, however the final result must be the sum of ”kinetic” deuterons and those found by (s)MST or formed via coalescence. The ”sMST + kinetic” sum is describes the experimental data shape quite well, however, underestimates it. There is no reason to make that similar sum for the probabilistic coalescence as it already overestimate the data.
The coalescence results (solid black line) differs from those from [11], but this can be explained by the smaller coalescence parameters $\Delta r=3.575$ fm and $\Delta p=0.285$ GeV/c used in previous study, which led to the smaller yield of deuterons.
@@
Figure 6: Rapidity density distribution of deuterons. Black solid lines are coalescence deuterons. The orange dash-dotted line are ”kinetic” deuterons (see text). The red dashed lines shows the sum of the ”sMST” and ”kinetic” deuterons. Black circles – NA49 experimental data (measured and mirrored) [2].
@@
Figure 7: Rapidity density distribution of $^{3}He$. Black solid lines are coalescence with spin-isospin coupling cut. The red dashed lines shows the ”sMST” results. Black circles – NA49 experimental data[2].
@@
Figure 9: Rapidity density distribution of deuterons (top panel) and $^{3}He$ (bottom) found by the coalescence based on the $E_{\text{bind}}<0$ cut. Deuteron lines are coupled with ”kinetic” deuterons. Black solid lines are for the initial ”M1” set of parameters, orange dash-dotted lines – for the ”M2” set, ”M3” set – red dashed lines and blue dotted lines are for the set ”M4”. Black circles – NA49 experimental data (measured and mirrored) [2].